# EMMOE: A Comprehensive Benchmark for Embodied Mobile Manipulation in Open Environments

### v1

@@ Abstract
Developing autonomous home robots controlled by natural language has long been a pursuit of human. While advancements in large language models (LLMs) and embodied intelligence make this goal closer, several challenges persist: the lack of a unified benchmark for more complex robot tasks, limited evaluation methods and metrics, data incompatibility between LLMs and mobile manipulation trajectories. To address these issues, we introduce Embodied Mobile Manipulation in Open Environments (EMMOE), which requires agents to interpret user instructions and execute long-horizon everyday tasks in continuous space. EMMOE seamlessly integrates high-level and low-level embodied tasks into a unified framework, along with three new metrics for more diverse assessment. Additionally, we collect EMMOE-100, which features in various task attributes, detailed process annotations, re-plans after failures, and two sub-datasets for LLM training. Furthermore, we design HomieBot, a sophisticated agent system consists of LLM with Direct Preference Optimization (DPO), light weighted navigation and manipulation models, and multiple error detection mechanisms. Finally, we demonstrate HomieBot’s performance and the evaluation of different models and policies.
Embodied AI, Benchmark, Mobile Manipulation
@@ 1 Introduction
Figure 1: Data example in EMMOE-100 dataset. A key feature of EMMOE-100 is the emphasis on the reasoning process and interleaved execution. In the shown task, the agent must check the fridge first. Otherwise, even if the agent finally gets a banana from the kitchen, it will not be considered as a success.
@@
Developing autonomous robots capable of exploring environments and performing various daily tasks through a single instruction has long been a pursuit in robotics. To achieve this goal, robots need to understand natural language instructions, make feasible plans, perceive and interact with dynamic environments, and equip with powerful navigation and manipulation skills. Traditional robot learning methods like imitation learning (IL) (Ho & Ermon, 2016) and reinforcement learning (RL) (Sutton, 2018) primarily focus on task-specific policies, but are always limited to short-horizon tasks and struggle to generalize to new tasks. Task and Motion Planning (TAMP) treats long-horizon mobile manipulation tasks as hybrid discrete-continuous search problems (Garrett et al., 2021) and addresses with a hierarchical architecture (Kaelbling & Lozano-Pérez, 2011): High-level task planning in discrete task space, low-level motion planning in continuous action space, and interleaved execution between two layers. However, the scope of TAMP remains limited. Despite various extensions (Garrett et al., 2020; Ren et al., 2024b; Chen et al., 2024), it still requires specific goal states and detailed scene configurations. The complexity and dynamism of real-world environments, and vague user descriptions make it highly challenging to meet these requirements.
@@
In recent years, the rapid development of LLM (Achiam et al., 2023; DeepSeek-AI et al., 2025) and embodied intelligence (Brohan et al., 2023b; Driess et al., 2023) has made this pursuit possible. The scope of each layer in TAMP has been largely broadened, and spawns various embodied tasks driven by language and vision. In high-level embodied tasks (Wu et al., 2023; Li et al., 2024a), LLMs have shown exceptional performance and powerful generalization capabilities. Advanced prompting techniques like Chain-of-Thought (COT) (Wei et al., 2022) have further enhanced the logical reasoning abilities of LLMs. Visual Language Models (VLMs) (Radford et al., 2021) enable agents to process visual inputs and understand current environments. Large Multimodal Models (LMMs) (Liu et al., 2024) extend the application of embodied agents to real-world scenarios. The most recent world models (Matsuo et al., 2022) and spatial models (Huang et al., 2024b) allow agents to more accurately perceive scene information and spatial relationships. In low-level embodied tasks, the emphasis of models has gradually shifted from single skill with specific objects (Shafiullah et al., 2023) to single skill with open-vocabulary objects (Fang et al., 2023b), then further to general models (Black et al., 2024), such as Visual Language Navigation (VLN) (Zhang et al., 2024e) and Visual Language Action (VLA) (Brohan et al., 2023a) models.
@@
However, several problems remain unresolved: 1) Lack of a comprehensive task and benchmark. Although significant progress has been made in various embodied tasks, there is still a gap between the current tasks and the envisioned language-driven intelligent robots. Meanwhile, existing embodied tasks always operate in isolation, neglecting the mutual influence caused by interleaved task execution. All these factors indicate the need for a more extensive task framework. By integrating different high-level and low-level embodied tasks, robots can achieve more advanced capabilities while enabling a unified evaluation of various embodied tasks. Each layer will constrain and influence the others, work collaboratively to accomplish the final task. 2) Inadequate evaluation methods and metrics. Embodied task planning involves causal dependencies between each step, and solutions are not absolute, thus making evaluations based solely on individual steps or the final state insufficient. Additionally, current evaluation methods rely heavily on simulators or PDDL files, which also limits the real-world deployment and application of agents. Furthermore, how to make more fine-grained evaluations of the entire agent system remains a challenge. 3) LLM grounding problems. Although LLMs excel in commonsense reasoning, they need to be grounded in current environments to produce realistic and practical outputs. Furthermore, due to the uncertainties and dynamic changes in the real world, LLMs must be able to make timely adjustments based on real-time feedback. However, the incompatibility between the conversation data required for LLMs and the trajectory data required for robotics increases the difficulty of grounding.
@@
To advance the development of intelligent autonomous robots, we propose EMMOE as an open challenge, which requires agents interpret user instructions and execute long-horizon everyday tasks in continuous space. Besides, we manually collect EMMOE-100, the first daily task dataset featuring various task attributes, detailed process annotations, analyses of each output, re-plans after failures. We also build Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) (Rafailov et al., 2024) sub-datasets to facilitate the alignment of LMM capabilities with specific embodied tasks. Finally, we introduce HomieBot, a sophisticated agent system that integrates both high-level and low-level models, as well as multiple error detection and adaptation mechanisms. An example of EMMOE challenge and EMMOE-100 tasks is shown in Fig.1.
@@ 2.2 EMMOE-100 Dataset
Leveraging Fetch robots (Fetch Robotics, 2020) and interactive environments from Habitat-Lab 2.0 (Szot et al., 2021), we collect EMMOE-100, a dataset consists of 100 complex everyday tasks. We sample 100 different episodes from Replica Challenge (Szot et al., 2021) to build simulation environments. In each scene, we design a daily mobile manipulation task and collect data. We manually control a Fetch robot to complete all tasks in continuous space and decompose execution trajectories into discrete subtasks, finally we obtain 966 subtasks in total. Each subtask consists of a pre-defined action, a target, and a low-level model selection. In addition to execution descriptions, they are also annotated with four first-person view images and detailed reasoning processes. Moreover, we intentionally design some failed subtasks and provide corresponding re-plans to enhance dataset robustness. To address grounding problems, we also construct SFT and DPO sub-datasets based on EMMOE-100, which will be formally introduced in Section 4.1.
To enhance task diversity and better align with human demands, we design tasks with five different attributes: Short-horizon tasks like pick something and place it somewhere. Long-horizon tasks which consist of at least ten subtasks. Open-ended tasks that allow multiple possible results and solutions. Logical tasks that provide vague description and require logical reasoning. Human-style tasks which are described in a natural conversation style. One task can possess multiple attributes simultaneously, as some of these attributes are not contradictory. Table 1 shows detailed comparisons with other mobile manipulation and embodied task datasets. We also provide detailed task statistics and a complete task list and in Appendix B.
@@ 2.3 Evaluation Metrics
The most fundamental metrics in embodied task planning are Success Rate (SR) and Goal-Condition Success (GC) (Shridhar et al., 2020). SR measures the proportion of successful trajectories, while GC is the ratio of goal conditions achieved at the end of a trajectory, a trajectory is considered successful only if GC reaches 100%. However, GC focuses only on the final result and relies on pre-defined state goals, thus failing to meet the requirements of our EMMOE tasks, which require fine-grained and language-based evaluations. Although some studies (Li et al., 2024a) conduct more fine-grained evaluations, they overlook the flexibility and coherence in embodied task planning and still rely on abstract terms. The success of an individual step may not contribute to the final success, and an output that differs from the ground truth but can complete the task in an alternative way should not be considered incorrect. Furthermore, fine-grained evaluation of the entire agent system remains a challenge. To overcome these limitations and provide more diverse evaluations, we propose the following new metrics. All details about definitions, calculations, and examples can be found in Appendix C.
@@
$\text{count}_{t}(\text{replan})$ is the number of re-plans in trajectory $t$, other symbol definitions are same as SER. SRR represents the effectiveness of re-planning and adaptability of the agent. When SRR reaches 100%, it indicates that the agent can adapt to all failures then successfully complete the task.
@@ 3.1 Overview
In this section, we will introduce how HomieBot accomplishes EMMOE tasks. HomieBot employs a hierarchical framework with communication mechanisms for interleaved execution. High-Level Planning (HLP) deals with embodied decision making and planning adaptation, while Low-Level Execution (LLE) handles continuous execution and provides feedback to HLP. We will describe HLP in Section 3.2 and LLE in Section 3.3. An overview is shown in Fig.2.
@@ 3.2 High Level Planning (HLP)
A long trajectory will be decomposed into several subtasks, the agent must continuously interact with the environment and adjust plans based on real-time feedback to ensure generated subtasks are practical. We design elaborate input and output instructions to facilitate dynamic adjustments during execution. Video-LLaVA (Lin et al., 2023) is selected as our planner model $M$ and fine-tuned with SFT and DPO sub-datasets, which will be described in Section 4.1.
@@
To help the LMM better understand current situations, the input instruction $I$ is as follows:
@@
In the visual component, four first-person view images $o_{1\sim 4}$ correspond to four directions: front, left, back and right. In the textual component, system information $s$ and user task $T$ remain constant throughout the conversation, reminding the agent of its responsibility. Feedback $f$ indicates the status of the last execution and detailed error information if failed, it will also be used to update other parts in $I$. Inventory $inv$ reflects the items currently held by the agent, primarily to prevent the generation of illogical actions, it is updated based on both $f$ and the type of the last action. Execution history $h$ logs all previous subtasks and their results. Once receiving $f$, the last subtask and its result will be logged in $h$. Besides, to better align with real-world scenarios, we prohibit directly inputting background information into the LMM (e.g. raw object data, Bird’s Eye View images etc.). The planner must explore the environment and enhance its intrinsic capabilities to generate more reasonable outputs.
@@
$A$ represents the analysis of each outputs, which is inspired by works like CoT (Wei et al., 2022). Before generating final outputs, planner model $M$ is expected to summarize previous executions and current situations, analyze what to do next, and propose the subsequent subtask $S$. To ensure the feasibility of the output, action in each subtask must be chosen from the available action list. Similarly, $m$ which represents the selected low-level models or policies, is also restricted to a given model list. target can be either an object or a spot, it should be observable in the provided images and deemed necessary to complete the task.
@@ 3.3 Low Level Execution (LLE)
LLE will convert $S$, $m$ and $inv$ from HLP into precise model-calling instructions. Error detection will be applied at different stages to monitor the execution process. Once the execution is completed or failed, environmental images and feedback will be sent back to HLP. We set up six skills based on the support of the simulator(see Table E1 for more details). Since the required information varies from models and would significantly impact the model performance, we establish two distinct settings to ensure fairness.
@@
More specifically, execution with background information means that the selected model needs precise position and state information of the target. As M3 (Gu et al., 2022) shows exceptional performance in all predefined skills when utilizing background information in Habitat, we choose it as the unique model choice $m$ in this setting. To ensure that M3 can recognize the natural language style $target$, we implement a name mapping for $target$ and adjust its original setting to better align with requirements of our tasks. In addition to text and image data, LLE also captures the execution process of each step and the entire trajectory data in video format. This means that HomieBot has the potential to bridge the gap between robot data and LMM data as the entire execution process is fully automated and annotated, users only need to set up the scene and input instructions. The video data can be utilized for IL in robotics, while the text and image data can be utilized for LMM training.
@@
Without background information means that the agent can only rely on the information captured by its sensors and the intrinsic abilities of low-level models to complete the task. As shown in Table E2, we set two manipulation models and two navigation models. For manipulation, RT-1-X (Padalkar et al., 2023) is for $Pick$ and $Place$, while Octo (Team et al., 2024b) is for $Open$ and $Close$. For navigation, NoMaD (Sridhar et al., 2024) specializes in image navigation and is suitable when $target$ is a spot or large object. PixNav (Cai et al., 2024) excels in pixel-level and object navigation, making it ideal when $target$ is a detectable object.
As the deployment of robots in the real world always demands high real-time performance and is constrained by hardware limitations, we prefer to choose lightweight models rather than the currently popular VLA models. This approach helps prevent the system from becoming too burdensome. Additionally, breaking down long-horizon tasks into action primitives would also reduce the performance requirements of low-level models. Compared to general-purpose end-to-end models, specialized lightweight models can complete the action while reducing time costs.
@@
To facilitate communication with HLP and provide more detailed error information, we further classify common errors into four main types and several subtypes. Logical error L1: The agent’s hands are already full but still attempts to pick/open/close; L2: The agent holds nothing but attempts to put; L3: The agent attempts to pick/put the object in a closed container; L4: The agent attempts to interact with a non-interactive object. Distance error D1: The agent stands too far and is unable to reach the target; D2: The agent is too close to the target and its arm is hindered from properly extending during interaction. Format Error F1: The output action or model is not in the available list; F2: The output target does not exist in the current scene or can not be recognized by low-level models. Execution Error E1: The limited capabilities of the low-level models or policies cause the failure; E2: Failed execution may result in the inventory information being accidentally updated. Furthermore, since these errors would occur at different stages, we also conduct multiple phases of error detection during execution. More classification and detection details are given in Appendix E.

### v2

@@ Abstract
Developing autonomous home robots controlled by natural language has long been a pursuit of humanity. While advancements in large language models (LLMs) and embodied intelligence make this goal closer, several challenges persist: the lack of a unified benchmark for more complex robot tasks, limited evaluation methods and metrics, data incompatibility between LLMs and mobile manipulation trajectories. To address these issues, we propose Embodied Mobile Manipulation in Open Environments (EMMOE), a benchmark that requires agents to interpret user instructions and execute long-horizon everyday tasks in continuous space. EMMOE seamlessly integrates high-level and low-level embodied tasks into a unified framework, along with three new metrics for more diverse assessment. Additionally, we collect EMMOE-100, which features in various task attributes, detailed process annotations, re-plans after failures, and two sub-datasets for LLM training. Furthermore, we design HomieBot, a sophisticated agent system consists of LLM with Direct Preference Optimization (DPO), light weighted navigation and manipulation models, and multiple error detection mechanisms. Finally, we demonstrate HomieBot’s performance and evaluations of different models and policies. All datasets, models, and codes can be found in our Project Website.
@@ 1 Introduction
Figure 1: Data example in EMMOE-100. A key feature of EMMOE-100 is the emphasis on the reasoning process and interleaved execution. In the shown task, the agent must check the fridge first. Otherwise, even if the agent finally gets a banana in the kitchen, it will not be considered as a success.
@@
Developing autonomous robots capable of performing various daily tasks through a single instruction has been a long-standing goal. To achieve this goal, robots need to understand natural language instructions, make feasible plans, perceive and interact with dynamic environments, and equip with powerful navigation and manipulation skills. Typical methods like imitation learning (IL) [1] and reinforcement learning (RL) [2] primarily focus on task-specific policies, but are always limited to short-horizon tasks and struggle to generalize to new tasks. Task and Motion Planning (TAMP) treats long-horizon mobile manipulation tasks as hybrid discrete-continuous search problems [3] and addresses with a hierarchical architecture [4]: High-level task planning in discrete task space, low-level motion planning in continuous action space, and interleaved execution between two layers. However, the scope of TAMP remains limited. Despite various extensions [5, 6, 7], it still requires specific goal states and detailed scene configurations. The complexity and dynamism of real-world environments, and vague user descriptions make it highly challenging to meet these requirements.
@@
In recent years, the rapid development of LLM [8, 9] and embodied intelligence [10, 11] has made this pursuit possible. The scope of each layer in TAMP has been largely broadened and spawns various embodied tasks driven by language and vision. In high-level embodied tasks [12, 13], LLMs have shown exceptional performance and powerful generalization capabilities. Advanced prompting techniques like Chain-of-Thought (COT) [14] have further enhanced the logical reasoning abilities of LLMs. Visual Language Models (VLMs) [15] enable agents to process visual inputs and understand current environments. Large Multi-modal Models (LMMs) [16] extend the application of embodied agents to real-world scenarios. The most recent world models [17] and spatial models [18] allow agents to more accurately perceive scene information and spatial relationships. In low-level embodied tasks, the emphasis of models has gradually shifted from single skill with specific objects [19] to single skill with open-vocabulary objects [20], then further to general models [21], such as Visual Language Navigation (VLN) [22] and Visual Language Action (VLA) [23] models.
@@
However, several problems remain unresolved: 1) Lack of a comprehensive task and benchmark. Although significant progress has been made in various embodied tasks, there is still a gap between the current tasks and the envisioned language-driven intelligent robots. Meanwhile, existing embodied tasks always operate in isolation, neglecting the mutual influence caused by interleaved task execution. By integrating different high-level and low-level embodied tasks, robots can achieve more advanced capabilities while enabling a unified evaluation of various embodied tasks. Each layer will constrain and influence the others, working collaboratively to accomplish the final task. 2) Inadequate evaluation methods and metrics. Embodied task planning involves causal dependencies between each step, and solutions are not absolute, thus making evaluations based solely on individual steps or the final state insufficient. Additionally, current evaluation methods rely heavily on simulators or PDDL files, which also limits the real-world deployment and application of agents. Furthermore, how to make more fine-grained evaluations of the entire agent system remains a challenge. 3) LLM grounding problems. Although LLMs excel in commonsense reasoning, they need to be grounded in current environments to produce realistic and practical outputs. Furthermore, due to the uncertainties and dynamic changes in the real world, LLMs must be able to make timely adjustments based on real-time feedback. However, the incompatibility between the conversation data required for LLMs and the trajectory data required for robotics increases the difficulty of grounding.
@@
To advance the development of intelligent autonomous robots, we propose EMMOE as an open challenge, which requires agents to interpret user instructions and execute long-horizon everyday tasks in continuous space. Besides, we manually collect EMMOE-100, the first daily task dataset featuring various task attributes, detailed process annotations, analyses of each output, re-plans after failures. We also build Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) [24] sub-datasets to facilitate the alignment of LMM capabilities with specific embodied tasks. Finally, we introduce HomieBot, a sophisticated agent system that integrates both high-level and low-level models, as well as multiple error detection and adaptation mechanisms. An example of EMMOE challenge and EMMOE-100 tasks is shown in Fig.1.
@@ 2.2 EMMOE-100 Dataset
By controlling Fetch Robots [25] in Habitat-Lab 2.0 [26], we collect EMMOE-100, a dataset consists of 100 complex everyday tasks. We sample 100 different scenarios from Replica Challenge [26] to build simulation environments. In each scene, we’ll first design a daily mobile manipulation task, then manually control a Fetch robot to complete the task in continuous space and decompose execution trajectories into discrete subtasks. Each subtask consists of an executable action, a target, and a low-level model selection, finally we obtain 966 subtasks in total. We also annotate each subtask with four first-person view images and detailed reasoning processes. Moreover, we intentionally design some failed subtasks and provide re-plans to enhance dataset robustness. To alleviate grounding problems, we construct SFT and DPO sub-datasets, which will be introduced in Section 4.1.
To enhance task diversity and better align with human demands, we design tasks with five different attributes: Short-horizon tasks like pick something and place it somewhere. Long-horizon tasks which consist of at least ten subtasks. Open-ended tasks that allow multiple results and solutions. Logical tasks that provide vague descriptions and require logical reasoning. Human-style tasks are described in a natural conversation style. One task can possess multiple attributes as some of these attributes are not contradictory. Table 1 shows detailed comparisons with other mobile manipulation and embodied task datasets. We also provide detailed task statistics in Appendix B.
@@ 2.3 Evaluation Metrics
The most fundamental metrics in embodied task planning are Success Rate (SR) and Goal-Condition Success (GC) [27]. SR measures the proportion of successful trajectories, while GC is the ratio of goal conditions achieved at the end of a trajectory. A trajectory is considered successful only if GC reaches 100%. However, GC focuses only on the final result and relies on pre-defined state goals, thus failing to meet the requirements of our EMMOE tasks, which require fine-grained and language-based evaluations. Although some studies [13] conduct more fine-grained evaluations, they overlook the flexibility and coherence in embodied task planning and still rely on abstract terms. The success of an individual step may not contribute to the final success, and an output that differs from the ground truth but can complete the task in an alternative way should not be considered incorrect. Furthermore, fine-grained evaluation of the entire agent system remains a challenge. To overcome these limitations and provide more diverse evaluations, we propose the following new metrics. All details about definitions and visible calculation examples can be found in Appendix C.
@@
$\text{count}_{t}(\text{replan})$ is the number of re-plans in trajectory $t$, other symbol definitions are the same as SER. SRR represents the effectiveness of re-planning and adaptability of the agent. When SRR reaches 100%, it indicates that the agent can adapt to all failures and then successfully complete the task.
@@ 3.1 Overview
In this section, we will introduce how HomieBot accomplishes EMMOE tasks. HomieBot employs a hierarchical framework with communication mechanisms for interleaved execution. High-Level Planning (HLP) deals with embodied decision making and planning adaptation, while Low-Level Execution (LLE) translates subtasks into continuous low-level controls and provides feedback to HLP. We will describe HLP in Section 3.2 and LLE in Section 3.3. A system overview is shown in Fig.2.
@@ 3.2 High Level Planning (HLP)
A long trajectory will be decomposed into several subtasks, the agent must continuously interact with the environment and adjust plans based on real-time feedback to ensure generated subtasks are practical. We design elaborate input and output instructions to facilitate dynamic adjustments during execution. Video-LLaVA [28] is selected as our planner model $M$ and fine-tuned with SFT and DPO sub-datasets, which will be described in Section 4.1.
@@
To help the LMM better understand current situations, the format of the input instruction $I$ is as follows:
@@
In the visual component, four first-person view images $o_{1\sim 4}$ correspond to four directions respectively. In the textual component, system information $s$ and user task $T$ remain constant throughout the conversation, reminding the agent of its responsibility. Feedback $f$ indicates the status of the last execution and detailed error information if failed, it will also be used to update other parts in $I$. Inventory $inv$ reflects the items currently held by the agent, primarily to prevent the generation of illogical actions, it is updated based on both $f$ and the type of the last action. Execution history $h$ logs all previous subtasks and their results. Once receiving $f$, the last subtask and its result will be logged in $h$. Besides, to better align with real-world scenarios, we prohibit directly inputting background information into the LMM (e.g. raw object data, Bird’s Eye View images etc.). The planner must explore the environment and enhance its intrinsic capabilities to generate more reasonable outputs.
@@
$A$ represents the analysis of each output, which is inspired by works like CoT [14]. Before generating final outputs, planner model $M$ is expected to summarize previous executions and current situations, analyze what to do next, and propose the subsequent subtask $S$. To ensure the feasibility of the output, action can only be chosen from the available action list. Similarly, $m$ which represents the selected low-level models or policies, is also restricted to a given model list. target can be either an object or a spot, it should be observable in the provided images and deemed necessary to complete the task.
@@ 3.3 Low Level Execution (LLE)
LLE will convert $S$, $m$ and $inv$ from HLP into precise model-calling instructions. Error detection will be applied at different stages to monitor the whole process. Once the execution is completed or failed, environmental images and feedback will be sent back to HLP. We set up six skills based on the support of the simulator(see Table E1). Since the required information varies from models and would significantly impact the model performance, we establish two distinct settings to ensure fairness.
@@
More specifically, execution with background information means that the selected model needs precise position and state information of the target. As M3 [29] shows exceptional performance in all skills when utilizing background information in Habitat, we choose it as the unique model choice $m$ in this setting. To make M3 adapt to our task requirements, we implement a name mapping for $target$ and adjust its original setting to better align with requirements of our tasks. In addition to text and image data, LLE also captures the execution process of each step and the entire trajectory data in video format. This means that HomieBot has the potential to bridge the gap between robot data and LMM data as the entire execution process is fully automated and annotated, users only need to set up the scene and input instructions. The video data can be utilized for IL in robotics, while the text and image data can be utilized for LMM training.
@@
Without background information means that the agent can only rely on the information captured by its sensors and the intrinsic abilities of low-level models to complete the task. As shown in Table E2, we set two manipulation models and two navigation models. For manipulation, RT-1-X [30] is used for $Pick$ and $Place$, while Octo [31] is set for $Open$ and $Close$. For navigation, NoMaD [32] specializes in image navigation and is suitable when $target$ is a spot or large object. PixNav [33] excels in pixel-level and object navigation, making it ideal when $target$ is a detectable object. As the deployment of robots in the real world always demands high real-time performance and is constrained by hardware limitations, we prefer to choose lightweight models rather than the currently popular VLA models to prevent the system from becoming too burdensome. Additionally, breaking down long-horizon tasks into action primitives would also reduce the performance requirements of low-level models. Compared to general-purpose end-to-end models, specialized lightweight models can complete the action while reducing time costs.
@@
To facilitate communication with HLP and provide more detailed error information, we further classify common errors into four main types and several subtypes. Logical error L1: The agent’s hands are already full but still attempts to pick/open/close; L2: The agent holds nothing but attempts to put; L3: The agent attempts to pick/put the object in a closed container; L4: The agent attempts to interact with a non-interactive object. Distance error D1: The agent stands too far and is unable to reach the target; D2: The agent is too close to the target and its arm is hindered from properly extending during interaction. Format Error F1: The output action or model is not in the available list; F2: The output target does not exist in the current scene or can not be recognized by low-level models. Execution Error E1: The limited capabilities of the low-level models or policies cause the failure; E2: Failed execution may result in the inventory information being accidentally updated. Furthermore, we conduct multiple phases of error detection during the whole process to guarantee the executions. More classification and detection details are given in Appendix E.